Consider a simple economy that produces only pies. The following table contains information on the economys money supply, velocity of money, price level, and output. For example, in 2015, the money supply was 100, the price of a pie was $4.00, and the economy produced 500 pies Fill in the missing values in the following table, rounding to the nearest cent when necessary Money Supply (Dollars) 100 103 Price Level (Dollars) 4.00 Quantity of Output (Pies) 500 500 Nominal GDP Velocity of Money (Dollars) Year 2015 2016 20 The money supply grew at a rate of % from 2015 to 2016, Since pie output did not change from 2015 to 2016 and the velocity of money ,the change in the money supply was reflected in changes in the price level. The inflation rate from 2015 to 2016 was
Answer #1
Year Money Supply $ Velocity of money Price Level Quantity Nominal GDP
2015 $ 100 20 $ 4.00 500 $ 2,000
2016 $ 103 20 $ 4.12 500 $ 2,060

Velocity of money, V = ?

MV = PQ

PO

V2015 = $$1000 = 20 15$100

Using the same formula we can determine the price level

large P = rac{MV}{Q}

$103*20 4.12 2016500

Nominal GDP = Price per unit * Quantity

Nominal GDP (2015) = $4 *500 = $ 2,000

Nominal GDP(2016) = $4.12*500 = $ 2,060

CP12016- CP12015 flation - ImtattoCPI2015

Inflation 2.000 * 100 = 3% 2,000

The money supply grew at a rate of 3% from 2015 to 2016. Since, the output of Pie did not change from 2015 to 2016 and the velocity of money remained the same , the change in money supply was reflected entirely in the price level. The inflation rate from 2015 to 2016 was 3%.

Please contact if having any query thank you.